An inductor has an inductance of 0.036 H. The voltage across this inductor is 41 V and has a frequency of 570 Hz. What is the current in the inductor?
Answer #1

Inductive reactance, Xl = 2 pi x 570 x 0.036 = 128.93 ohm

Current, I = V /Xl = 41/128.93 = 0.32 A

Comment in case any doubt please rate my answer ....